OpenStudy (anonymous):

A watermelon composed of 99% of water, weighted 10kg. After a while, some of water evaporated, leaving the watermelon composed of 98% of water. What is the new weight of water-melon?

OpenStudy (anonymous):

is % - x - of 100

OpenStudy (anonymous):

then cross multiply

OpenStudy (anonymous):

sup homies the answer is

OpenStudy (anonymous):

99% OF 10kg ....

OpenStudy (anonymous):

This is a tricky question...

OpenStudy (anonymous):

x 99 - x - 10 100

OpenStudy (anonymous):

Would that mean- 100x=990?

OpenStudy (anonymous):

yea then divide

OpenStudy (anonymous):

9.9?

OpenStudy (espex):

While rounded the answers are the same, I would offer a slight modification from @mads4566 approach. Use the ratio of water to weight as: \[\frac{.99}{10} = \frac{.98}{x}\] And cross multiply as suggested.

OpenStudy (anonymous):

Thanks @eSpeX

OpenStudy (espex):

You're welcome, @mads4566 got the ball rolling too. :)

OpenStudy (anonymous):

ahah yeah (: Thank you BOTH

OpenStudy (anonymous):

a cute question can also solve like this if the ten pounds of watermelon starts at 99% water, then you have 9.9 pounds of water and .1 pound of solid then some water evaporates, but you still have .1 pound of solid, which is 2% of the total solve 2% of what number is .1 via \(.1\div .02\) do get the answer

OpenStudy (anonymous):

Sorry but im confused now...

OpenStudy (anonymous):

i am not sure what the first answers above give, but it looks like they give the wrong answer \[\frac{.1}{.02}=\frac{10}{2}=5\] so the weight is 5 pounds

OpenStudy (anonymous):

it is a tricky question, so lets go slow

OpenStudy (espex):

Are you saying that a .01 loss of moisture equates to a 5kg loss?

OpenStudy (anonymous):

you have 10 pounds that is water and solid, 99% water so 1% solid and 1% of 10 is 0.1

OpenStudy (anonymous):

so you have .1 pound of solid, that that does not evaporate now a day later you have 98% water, meaning the rest , the other 2% is the solid

OpenStudy (anonymous):

so the question is now this: 0.1 pounds is 2% of what number?

OpenStudy (anonymous):

and the answer is, 0.1 is 2% of 5

OpenStudy (anonymous):

@eSpeX yes, that is what i am saying

OpenStudy (espex):

@satellite73 It is not a 2% loss of total weight, it is a 1% loss of moisture.

OpenStudy (anonymous):

ok we start again lets change the probem slightly to make it easier suppose instead of 10 pound you have 100 pounds of watermelon, and it is 99% water, hence 99 pounds water, and one pound solid

OpenStudy (espex):

By your calculations, assuming I am reading it right, a 1% loss of moisture resulted in 1/2 of the total weight evaporating.

OpenStudy (anonymous):

now after a day, there is still one pound of solid, but the mixture is now 98% water and therefore 2% not water

OpenStudy (anonymous):

so you have 1 pound that is 2% of the total, and 1 is 2% of 50

OpenStudy (anonymous):

i was stil right with the is/of x %/100 right?

OpenStudy (anonymous):

so a day later you have 50 pounds that would be 49 pounds of water and 1 pound of solid

OpenStudy (anonymous):

and you can easily check that the answer is right, because if i have 50 pounds, 49 of it water and 1 pound solid, then i have 98% water and 1 pound solid the answer is correct

OpenStudy (anonymous):

i think we lost @Genuine, but the 9.9 is the obvious but totally wrong answer

OpenStudy (anonymous):

oh well, looks like genuine is gone. alas

OpenStudy (espex):

I see where you're going with this, it is the part that 98% now makes up only 1/2 of what 99% did that I am struggling with.

OpenStudy (anonymous):

9.9 kg